OpenStudy (anonymous):

how to find the sqrt of 12.57

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!}\)

OpenStudy (anonymous):

without a calculator?

OpenStudy (unklerhaukus):

taylor series?

OpenStudy (anonymous):

Assuming you mean with a calculator, just use the square root function: \[\sqrt{12.57}\] = 3.54541958 = 3.55

OpenStudy (anonymous):

wow unkle is typing alot

OpenStudy (unklerhaukus):

\[ \sqrt{12.57}\\ =(16-3.43)^{1/2}\\ =4\left(1-\frac{3.43}{16}\right)^{1/2}\\ =4\left[1+\tfrac12\Big(\frac{-3.43}{16}\Big)+\frac{\tfrac12(\tfrac12-1)}{2!}\Big(\frac{-3.43}{16}\Big)^2+\mathcal O\big((\tfrac{-3.43}{16})^3\big)\right]\\ =4\left[1+\Big(\frac{-3.43}{32}\Big)-\frac{\tfrac14}{2}\Big(\frac{-3.43}{16}\Big)^2+\mathcal O\big((\tfrac{-3.43}{16})^3\big)\right] \\\approx4\times\left(1-\frac{3.43}{32}-\tfrac18\times\left(\frac{3.43}{16}\right)^2\right) \\=3.548\]

OpenStudy (anonymous):

holy mcdonald!!!

OpenStudy (unklerhaukus):

close enough

OpenStudy (anonymous):

uncle from sqrt(12.57) where did you get (16 - 3.43)^1/2

OpenStudy (unklerhaukus):

well i know the square root of 16,

OpenStudy (anonymous):

isnt it 4

OpenStudy (anonymous):

16-3.43 = 12.57. He did it as he knew the sqrt of 16.

OpenStudy (unklerhaukus):

yes

OpenStudy (anonymous):

ohhh

OpenStudy (unklerhaukus):

then factor out the four , and apply a binomial series

OpenStudy (dls):

Alternate way: Let \(\Large f(x) = \sqrt x\) We need for x=12.57 As we now.... \[\Large f(x+ \Delta x) = f(x) + f' (x) [ \Delta x]\] \[\Large f(16-3.43) = f(16) + \frac{1}{2 \sqrt 16 } \times ( - 3.43)\] \[\Large f(12.57)= 4- \frac{3.43}{8} = > \frac{32- 3.43}{8} = 3.57\] Note : this is a rough approximation.

OpenStudy (anonymous):

that earns a metal for hard work

OpenStudy (unklerhaukus):

thats kinda the same method